# Vesterbacka Rekathati Kurtz Sikora Toftgård National Library of SwedenSweden Swedish Whispers; Leveraging a Massive Speech Corpus for Swedish Speech Recognition

### v1

@@ Abstract
This work presents a suite of fine-tuned Whisper models for Swedish, trained on a dataset of unprecedented size and variability for this mid-resourced language. As languages of smaller sizes are often underrepresented in multilingual training datasets, substantial improvements in performance can be achieved by fine-tuning existing multilingual models, as shown in this work. This work reports an overall improvement across model sizes compared to OpenAI’s Whisper evaluated on Swedish. Most notably, we report an average 47% reduction in WER comparing our best performing model to OpenAI’s whisper-large-v3, in evaluations across FLEURS, Common Voice, and NST.
@@
The abundance of openly available audio data in English enables pre-training and fine-tuning of speech-to-text models on hundreds of thousands to millions of hours of recorded speech. As a result, automatic speech recognition (ASR) systems are approaching human level robustness in English. Other languages’ performance in multilingual speech recognition tends to stand in proportion to the amount of data included from the language – or the language family – in question. For low- to mid-resource languages with fewer speakers, such as Swedish, the amount of openly available data may be limited, and as a consequence these languages tend to be underrepresented in large-scale efforts to train multilingual speech recognition systems. This results in suboptimal performance of ASR applications, such as voice-based customer service, and automatic captioning. Language models and ASR models are often trained on massively multilingual datasets. By continuing training these multilingual models for a specific language, the performance is greatly improved compared to the performance of multilingual models. This is necessary for low- and mid-resource languages to remain competitive and sovereign in the rapid AI development of today. Similarly to previous work for Norwegian [1], we fine-tune Whisper for all canonical sizes on our massive dataset for Swedish.
@@ 3 A massive Swedish speech corpus
OpenAI’s whisper-large-v3222https://huggingface.co/openai/whisper-large-v3 has been trained on of multilingual speech, of which are in Swedish. While [3] report that all increases in dataset size lead to improved performance, the improvement slightly stagnates after a dataset size of around . With these numbers in mind, a dataset consisting of of transcribed Swedish was constructed.
@@
In this work, KBLab at the National Library of Sweden has collaborated with the Riksdag administration (i.e. the parliament administration in Sweden), Sweden’s Television (SVT), and the Institute for Language and Folklore (Institutet för språk och folkminnen), and together we have been able to construct a transcribed Swedish dataset of unprecedented size and variability. These dataset types can be split into four categories, which are presented below.
@@ 3.1 Subtitles
In our filtering, we select only Swedish TV broadcasts and web content where metadata indicates the subtitle track is in Swedish. To ensure the speech is also in Swedish, we perform language detection using OpenAI’s whisper-large-v3. The final subtitle dataset is thus constructed out of audio identified as Swedish paired with Swedish subtitles.
@@ 3.2 Parliamentary recordings
As a part of this project, we have been granted access to all digitized audio recordings of the Swedish Riksdag since 1966. These recordings were matched and force-aligned against protocol transcripts of the debates. The resulting speech corpus is named RixVox-v2, and consists of of transcribed speech recorded before 2003 and after 2003. Apart from its sheer size, this corpus also exhibits a wide variation in Swedish dialects as well as providing a both modern and historic resource for political spoken Swedish. The total amount of audio recordings used in this project is , of which almost pass the “Stage 1” quality filters outlined in Section 3.5. We publish the entire dataset RixVox-v2 on Hugging Face with a permissive license to further research efforts in Swedish speech technology.
@@ 3.4.2 Common Voice
Mozilla is collecting an open source, crowd-sourced, multilingual “dataset of voices” called Common Voice [7]. The dataset is the largest publicly available speech dataset, of which Swedish constitutes of recordings. The train, validation, and test splits of Common Voice 16.1 are all used as hold-out test sets.
@@ 3.5 Data preprocessing
Candidate chunks to be used during training are created by combining and packing our observations to chunks of up to length based on timestamp metadata. Shorter chunks are additionally sampled by setting a smaller maximum chunk length, in order to boost the model’s robustness to variable audio input lengths.
@@
Recall-Oriented Understudy for Gisting Evaluation score (ROUGE-N) [11] is a recall-based metric that measures the overlap of n-grams between a model’s output and a reference text.
@@
Stage 1 The first category is defined by a set of relaxed values for CER and BLEU, comparing the original transcription to that of whisper-large-v3 and VoxRex. Although the whisper-large-v3 transcriptions are generally of higher quality than those from VoxRex, both transcriptions are used to safeguard against potential hallucinations from Whisper. We apply dataset specific BLUE thresholds based on qualitative assessments, where the main aim during “Stage 1” is to filter out most of the low signal and pure noise observations where ground truth has very low or no overlap with the audio. CER thresholds comparing the first and last 10 characters of ground truth and machine transcriptions are used only to determine which samples to train with timestamp tokens during this stage. The hypothesis behind this category of data is that even though the transcriptions are not perfect, they are still good enough to learn from.
Stage 2 The second category, denoted “Stage 2”, is characterized by a set of higher requirements for the quality of the transcriptions, to ensure that only the most accurate examples from the training data are included in the second stage. If the previous category was characterized by the width of the variation of Swedish, this category will instead be characterized by the transcription style. Additionally, in Stage 2 filtering, we introduce another metric, ROUGE-N, to further ensure high transcription accuracy. During filtering, we calculate BLEU score and a weighted combined ROUGE-N score using weights that place higher emphasis on longer n-grams. We assign weights 0, 0.25, 0.5, 0.25, respectively to 1-, 2-, 3- and 4-grams. A CER threshold of 0.2 or lower for the first and last 10 characters is used here to determine whether to include the sample.
@@ 3.5.1 Non-speech and timestamps
A portion of the training data is prepared with timestamp tokens added at the subtitle block or sentence level, with the goal of retaining the original Whisper model’s ability to output timestamps. In order to provide a consistent signal to the model, we only train with timestamps on those examples where the CER between ground truth and the first and last 10 characters of an ASR transcription is equal to 0.2 or lower. Timestamps are sampled into the training at a rate of roughly 50% during “Stage 2”, and at lower single digit percentage rates in “Stage 1”.
@@ 5 Results and discussion
We report test results as a comparison between OpenAI’s Whisper and our models evaluated using the metrics word error rate (WER) and BLEU scores. The splits of three hold-out datasets which have not been employed for validation purposes during training are used: namely FLEURS (train and test set), NST (test set), and Common Voice (train, validation, and test set). Common Voice and FLEURS are not part of our training data and as such they represent the models’ out-of-domain performance. There is a significant improvement in performance for all model sizes trained in this paper compared to OpenAI’s Whisper, which can be seen in Table 3 for WER and Table 4 for BLEU.
@@
Most notable is the improvement in WER for the smaller model sizes, where the WER for the OpenAI Whisper-tiny model evaluated on FLEURS resulted in 59.2, a number which has rendered the model more or less inutile. We report an improvement in WER to 13.2 for the smallest model size (tiny). The second noteworthy observation is that our Whisper-small outperforms OpenAI’s Whisper-large as evaluated on these benchmarks. This result has a positive implication on the amount of compute needed for transcription tasks, as the same performance can be achieved for a much smaller model that in turn consumes significantly less compute during inference. Most notably, we report an average 47% reduction in WER comparing our best performing model to OpenAI’s Whisper large-v3, in evaluations across FLEURS, Common Voice, and NST.
@@ 5.1 Future work
The data preprocessing pipeline as outlined in Section 3.5 is relying on the transcriptions from OpenAI’s whisper-large-v3 to select high-quality training data. An obvious drawback with this approach is that one can not differentiate if the training data is of bad quality (e.g. not identified as Swedish) or if the model used to transcribe is not good enough. One risk is that this approach is rejecting training data with difficult dialects as the OpenAI model has not been trained on this type of speech before. By using our fine-tuned Whisper-large for the transcription step in the data preprocessing pipeline, one could potentially select more high-quality training data as well as training data with rare dialects. So data that had otherwise been filtered out due to a bad model performance, could in fact be included in the training data now, and further widen the range of quality in the training data. With the same reasoning, the use of a better Wav2Vec2.0 model than the VoxRex used in the data preprocessing pipeline would be beneficial.

### v2

@@ Abstract
This work presents a suite of fine-tuned Whisper models for Swedish, trained on a dataset of unprecedented size and variability for this mid-resourced language. As languages of smaller sizes are often underrepresented in multilingual training datasets, substantial improvements in performance can be achieved by fine-tuning existing multilingual models, as shown in this work. This work reports an overall improvement across model sizes compared to OpenAI's Whisper evaluated on Swedish. Most notably, we report an average 47% reduction in WER comparing our best performing model to OpenAI's whisper-large-v3, in evaluations across FLEURS, Common Voice, and NST.
@@
The abundance of openly available audio data in English enables pre-training and fine-tuning of speech-to-text models on hundreds of thousands to millions of hours of recorded speech. As a result, automatic speech recognition (ASR) systems are approaching human level robustness in English. Other languages' performance in multilingual speech recognition tends to stand in proportion to the amount of data included from the language – or the language family – in question. For low- to mid-resource languages with fewer speakers, such as Swedish, the amount of openly available data may be limited, and as a consequence these languages tend to be underrepresented in large-scale efforts to train multilingual speech recognition systems. This results in suboptimal performance of ASR applications, such as voice-based customer service, and automatic captioning. Language models and ASR models are often trained on massively multilingual datasets. By continuing training these multilingual models for a specific language, the performance is greatly improved compared to the performance of multilingual models. This is necessary for low- and mid-resource languages to remain competitive and sovereign in the rapid AI development of today. Similarly to previous work for Norwegian [1], we fine-tune Whisper for all canonical sizes on our massive dataset for Swedish.
@@ 3 A massive Swedish speech corpus
OpenAI's whisper-large-v3222https://huggingface.co/openai/whisper-large-v3 has been trained on of multilingual speech, of which are in Swedish. While [3] report that all increases in dataset size lead to improved performance, the improvement slightly stagnates after a dataset size of around . With these numbers in mind, a dataset consisting of of transcribed Swedish was constructed.
@@
In this work, KBLab at the National Library of Sweden has collaborated with the Riksdag administration (i.e. the parliament administration in Sweden), Sweden's Television (SVT), and the Institute for Language and Folklore (Institutet för språk och folkminnen), and together we have been able to construct a transcribed Swedish dataset of unprecedented size and variability. These dataset types can be split into four categories, which are presented below.
@@ 3.1 Subtitles
In our filtering, we select only Swedish TV broadcasts and web content where metadata indicates the subtitle track is in Swedish. To ensure the speech is also in Swedish, we perform language detection using OpenAI's whisper-large-v3. The final subtitle dataset is thus constructed out of audio identified as Swedish paired with Swedish subtitles.
@@ 3.2 Parliamentary recordings
As a part of this project, we have been granted access to all digitized audio recordings of the Swedish Riksdag since 1966. These recordings were matched and force-aligned against protocol transcripts of the debates, a part of the research corpus, the Swedish Parliament Corpus[6]. The resulting speech corpus is named RixVox-v2, and consists of of transcribed speech recorded before 2003 and after 2003. Apart from its sheer size, this corpus also exhibits a wide variation in Swedish dialects as well as providing a both modern and historic resource for political spoken Swedish. The total amount of audio recordings used in this project is , of which almost pass the ``Stage 1'' quality filters outlined in Section 3.5. We publish the entire dataset RixVox-v2 on Hugging Face with a permissive license to further research efforts in Swedish speech technology.
@@ 3.4.2 Common Voice
Mozilla is collecting an open source, crowd-sourced, multilingual ``dataset of voices'' called Common Voice [8]. The dataset is the largest publicly available speech dataset, of which Swedish constitutes of recordings. The train, validation, and test splits of Common Voice 16.1 are all used as hold-out test sets.
@@ 3.5 Data preprocessing
Candidate chunks to be used during training are created by combining and packing our observations to chunks of up to length based on timestamp metadata. Shorter chunks are additionally sampled by setting a smaller maximum chunk length, in order to boost the model's robustness to variable audio input lengths.
@@
Recall-Oriented Understudy for Gisting Evaluation score (ROUGE-N) [12] is a recall-based metric that measures the overlap of n-grams between a model's output and a reference text.
@@
Stage 1 The first category is defined by a set of relaxed values for CER and BLEU, comparing the original transcription to that of whisper-large-v3 and VoxRex. Although the whisper-large-v3 transcriptions are generally of higher quality than those from VoxRex, both transcriptions are used to safeguard against potential hallucinations from Whisper. We apply dataset specific BLUE thresholds based on qualitative assessments, where the main aim during ``Stage 1'' is to filter out most of the low signal and pure noise observations where ground truth has very low or no overlap with the audio. CER thresholds comparing the first and last 10 characters of ground truth and machine transcriptions are used only to determine which samples to train with timestamp tokens during this stage. The hypothesis behind this category of data is that even though the transcriptions are not perfect, they are still good enough to learn from.
Stage 2 The second category, denoted ``Stage 2'', is characterized by a set of higher requirements for the quality of the transcriptions, to ensure that only the most accurate examples from the training data are included in the second stage. If the previous category was characterized by the width of the variation of Swedish, this category will instead be characterized by the transcription style. Additionally, in Stage 2 filtering, we introduce another metric, ROUGE-N, to further ensure high transcription accuracy. During filtering, we calculate BLEU score and a weighted combined ROUGE-N score using weights that place higher emphasis on longer n-grams. We assign weights 0, 0.25, 0.5, 0.25, respectively to 1-, 2-, 3- and 4-grams. A CER threshold of 0.2 or lower for the first and last 10 characters is used here to determine whether to include the sample.
@@ 3.5.1 Non-speech and timestamps
A portion of the training data is prepared with timestamp tokens added at the subtitle block or sentence level, with the goal of retaining the original Whisper model's ability to output timestamps. In order to provide a consistent signal to the model, we only train with timestamps on those examples where the CER between ground truth and the first and last 10 characters of an ASR transcription is equal to 0.2 or lower. Timestamps are sampled into the training at a rate of roughly 50% during ``Stage 2'', and at lower single digit percentage rates in ``Stage 1''.
@@ 5 Results and discussion
We report test results as a comparison between OpenAI's Whisper and our models evaluated using the metrics word error rate (WER) and BLEU scores. The splits of three hold-out datasets which have not been employed for validation purposes during training are used: namely FLEURS (train and test set), NST (test set), and Common Voice (train, validation, and test set). Common Voice and FLEURS are not part of our training data and as such they represent the models' out-of-domain performance. There is a significant improvement in performance for all model sizes trained in this paper compared to OpenAI's Whisper, which can be seen in Table 3 for WER and Table 4 for BLEU.
@@
Most notable is the improvement in WER for the smaller model sizes, where the WER for the OpenAI Whisper-tiny model evaluated on FLEURS resulted in 59.2, a number which has rendered the model more or less inutile. We report an improvement in WER to 13.2 for the smallest model size (tiny). The second noteworthy observation is that our Whisper-small outperforms OpenAI's Whisper-large as evaluated on these benchmarks. This result has a positive implication on the amount of compute needed for transcription tasks, as the same performance can be achieved for a much smaller model that in turn consumes significantly less compute during inference. Most notably, we report an average 47% reduction in WER comparing our best performing model to OpenAI's Whisper large-v3, in evaluations across FLEURS, Common Voice, and NST.
@@ 5.1 Future work
The data preprocessing pipeline as outlined in Section 3.5 is relying on the transcriptions from OpenAI's whisper-large-v3 to select high-quality training data. An obvious drawback with this approach is that one can not differentiate if the training data is of bad quality (e.g. not identified as Swedish) or if the model used to transcribe is not good enough. One risk is that this approach is rejecting training data with difficult dialects as the OpenAI model has not been trained on this type of speech before. By using our fine-tuned Whisper-large for the transcription step in the data preprocessing pipeline, one could potentially select more high-quality training data as well as training data with rare dialects. So data that had otherwise been filtered out due to a bad model performance, could in fact be included in the training data now, and further widen the range of quality in the training data. With the same reasoning, the use of a better Wav2Vec2.0 model than the VoxRex used in the data preprocessing pipeline would be beneficial.